OpenStudy (anonymous):

HELP PLEASE? <3 You inherit $25000 and deposit it into an account that earns 4.5% annual interest compounded quarterly. How much money will be in the account after ten years?

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large A = P\left(1+\frac{r}{n}\right)^{n*t}\]

jimthompson5910 (jim_thompson5910):

In this case, P = 25000 r = 0.045 (the decimal form of 4.5%) n = 4 (compounding quarterly) t = 10

OpenStudy (anonymous):

I got 25000 (1+(0.045/4)^4(10)

jimthompson5910 (jim_thompson5910):

now evaluate

OpenStudy (anonymous):

When I calculate it comes out to 261441.2716. I'm not sure that seems right.

jimthompson5910 (jim_thompson5910):

that's way too big

jimthompson5910 (jim_thompson5910):

10 is in the exponent

jimthompson5910 (jim_thompson5910):

so it should be 25000*(1+0.045/4)^(4*10)

jimthompson5910 (jim_thompson5910):

evaluate that

OpenStudy (anonymous):

Okay, so I put it in wrong. So, it is 39109.42163, right?

jimthompson5910 (jim_thompson5910):

Very good, \[\Large A = P\left(1+\frac{r}{n}\right)^{n*t}\] \[\Large A = 25000\left(1+\frac{0.045}{4}\right)^{4*10}\] \[\Large A \approx 39,109.4216347769\] \[\Large A \approx 39,109.42\]

jimthompson5910 (jim_thompson5910):

So you'll have $39,109.42 in the account in 10 years

OpenStudy (anonymous):

Lovely~! Thank you. Would you mind if I ask you for help on one more part of the question set?

jimthompson5910 (jim_thompson5910):

Earning 39,109.42 - 25,000 = 14,109.42 in interest

jimthompson5910 (jim_thompson5910):

sure go ahead

OpenStudy (anonymous):

If the interest were to be compounded continuously at 4.5%, how much money would be in the account after 10 years? And how many years you would reach your goal of 65,000, if the interest is compounded continuously. I got 21.4 years for how long it takes to get 65000.

OpenStudy (anonymous):

And thank you for helping me with all of this. <3

jimthompson5910 (jim_thompson5910):

So you were able to answer the second part of this? but not the first part?

OpenStudy (anonymous):

I did it on my calculator quickly after you helped me just now.

jimthompson5910 (jim_thompson5910):

how did you do it on your calculator?

OpenStudy (anonymous):

When you showed me how to put in the formula, I changed the year, putting in 20. Since 10 gives you around 39 grand, So I looked to see if 20 would double it, but when it didn't I looked to se how much more I would have to put in and came up with around 21.

OpenStudy (anonymous):

Am I not supposed to do it like that? o.o

jimthompson5910 (jim_thompson5910):

well notice how it says "compounded continuously" so the formula is going to be (slightly) different

OpenStudy (anonymous):

May I ask how so?

jimthompson5910 (jim_thompson5910):

You will now use this formula \[\Large A = Pe^{rt}\] P is the amount you deposit or invest r is the interest rate (in decimal form) t is the time in years 'e' is a special constant (like pi) e = 2.71828182846...

jimthompson5910 (jim_thompson5910):

The explanation as to why you use this formula for compounding continuously is a bit lengthy and it involves some advanced theory Here is the explanation http://www-stat.wharton.upenn.edu/~waterman/Teaching/IntroMath99/Class04/Notes/node13.htm Unfortunately it requires some background in calculus to fully understand how they're deriving the formula

jimthompson5910 (jim_thompson5910):

anyways, you're taking the interest you've earned ($14,109.42) and you are depositing it into another account which continuously compounds interest for 10 years So P = 14,109.42 r = 0.045 t = 10 and you plug all that into \[\Large A = Pe^{rt}\]

OpenStudy (anonymous):

So, 14,109.42e^(0.045*10)?

OpenStudy (anonymous):

Doing it like that I get 39765.7485.

jimthompson5910 (jim_thompson5910):

what did you type in to get that? that's too high

OpenStudy (anonymous):

14,109.42x^(0.045*10). I can't find e on the calculator.

OpenStudy (anonymous):

Would the x change it that much?

jimthompson5910 (jim_thompson5910):

what kind of calculator do you have?

OpenStudy (anonymous):

A graphing calculator. I could use the alpha, but I'm not sure how much of a difference that would make to change it to a e.

jimthompson5910 (jim_thompson5910):

a texas instruments calculator?

OpenStudy (anonymous):

Changing it to a e in the equation made it 0. And yes. A TI-84 Plus.

jimthompson5910 (jim_thompson5910):

see the "LN" button (under the "Log" button)?

OpenStudy (anonymous):

Yes.

jimthompson5910 (jim_thompson5910):

above the LN is "e^x" so you hit the "2nd" key, then hit LN to get to e^x

OpenStudy (anonymous):

Okay~! So, it would be 22127.97532.

jimthompson5910 (jim_thompson5910):

good, that rounds to $22,127.98

OpenStudy (anonymous):

Ahhh...~! Okay! Thank you so much! ^^

jimthompson5910 (jim_thompson5910):

For the second part, the answer you got (21.4 years) is incorrect

OpenStudy (anonymous):

Is it?

jimthompson5910 (jim_thompson5910):

Yes, you need to solve for t in... A = P*e^(r*t) 65000 = 14109.42*e^(0.045*t)

OpenStudy (anonymous):

I got 0. o.o

jimthompson5910 (jim_thompson5910):

hmm odd

jimthompson5910 (jim_thompson5910):

how are you getting 0?

OpenStudy (anonymous):

Not sure. I'm putting it in the way you wrote it down. Could it be the letter?

jimthompson5910 (jim_thompson5910):

I'll show you how I solved for t. A = P*e^(r*t) 65000 = 14109.42*e^(0.045*t) 65000/14109.42 = e^(0.045*t) 4.60685 = e^(0.045*t) ln(4.60685) = ln(e^(0.045*t)) ln(4.60685) = 0.045*t*ln(e) ln(4.60685) = 0.045*t*1 ln(4.60685) = 0.045*t ln(4.60685)/0.045 = t t = ln(4.60685)/0.045 t = 33.945429

jimthompson5910 (jim_thompson5910):

the "ln" is really "LN" just with lowercase letters

jimthompson5910 (jim_thompson5910):

So it takes roughly 33.945429 years to go from $14,109.42 to $65,000

OpenStudy (anonymous):

That is so much. How do you retain it all? ;w;

jimthompson5910 (jim_thompson5910):

tons and tons of practice

OpenStudy (anonymous):

Oh, goodness. ><

jimthompson5910 (jim_thompson5910):

The first step after you plug everything in is to isolate e^(0.045*t)

jimthompson5910 (jim_thompson5910):

once you do that, you need to isolate t

jimthompson5910 (jim_thompson5910):

and you use natural logs (LN) to do so

OpenStudy (anonymous):

You explain this better than my teacher. o3o

jimthompson5910 (jim_thompson5910):

Thanks. I'm glad it's clicking now.

jimthompson5910 (jim_thompson5910):

Also, you'll use these log rules http://www.purplemath.com/modules/logrules.htm in this specific case, I used rule # 3

OpenStudy (anonymous):

Okay. Thank you. :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

Can I ask you a general question? ><;

OpenStudy (anonymous):

Just one more.

jimthompson5910 (jim_thompson5910):

Go ahead

OpenStudy (anonymous):

I am working on the next question right now and it asked me to identify the annual growth rate and the growth factor. What is the difference?

jimthompson5910 (jim_thompson5910):

What's the full problem you're working on?

OpenStudy (anonymous):

you deposit $2000 in an account that earns 5% annual interest compounded monthly.

OpenStudy (anonymous):

There are two parts before that and I am on the last part.

jimthompson5910 (jim_thompson5910):

how long are you depositing it for?

OpenStudy (anonymous):

The first question was; What will your balance be after 2 years? The second question; Estimate how long it would take for your investment to double. I'm not sure if that tells you. It only says annually in the third question; Identify the annual growth rate and the growth factor.

jimthompson5910 (jim_thompson5910):

it seems like this is a trick question because I think the annual growth rate is simply 5%

OpenStudy (anonymous):

How would you come across that if you don't mind me asking?

jimthompson5910 (jim_thompson5910):

it's given to you the 5% APR

OpenStudy (anonymous):

OH! Okay. I read over that. ^^;;

jimthompson5910 (jim_thompson5910):

as for the annual growth factor, that would be 1 + r/n = 1 + 0.05/12 = 1.0041667 which is approximate

OpenStudy (anonymous):

Okay~! That makes more sense. Thank you. <3

jimthompson5910 (jim_thompson5910):

you're welcome